# Penalizing Infeasible Actions and Reward Scaling in Reinforcement Learning with Offline Data

### v1

@@ 3.1 Extrapolation Error for OOD Actions Outside the Convex Hull of In-distribution Samples
To investigate this, we analyzed how the max-Q action, which maximizes the learned Q-function for a given state, differs from the dataset actions across various D4RL (Fu et al., 2020) datasets, similar to the approach of Yue et al. (2024). We trained $Q_{\phi}$ using observed state-action pairs from the dataset (offline SARSA; Sutton, 2018; Kumar et al., 2022) and then derived the max-Q policy using the loss function $\mathcal{L}_{\text{max-}Q}(\phi)=\mathbb{E}_{s\sim\mathcal{D}}\left[\text{max}_{a}Q_{\phi}(s,a)\right]$
@@ 4.1.1 Didactic Example
Figure 5: Results of training on a toy dataset using ReLU MLPs with vanilla regression (the ‘No LN’ column) and with LN (the ‘LN’ column), varying $c_{\text{reward}}$. The ‘LN (Wider Range)’ column shows results under the same conditions as ‘LN’, but with the $x_{1}$ and $x_{2}$ axes extended to the range $[-5,5]$. For more examples, please refer to Appendix D.
@@
As illustrated in the ‘No LN’ column, the Q-values exhibit catastrophic overestimation in $\mathcal{A}_{\text{OOD-out}}(s)$. Integrating LN helps mitigate the overestimation caused by linear extrapolation; however, $\mathcal{A}_{\text{OOD-out}}(s)$ remains overestimated relative to the in-distribution region. As the reward scale increases with LN applied (see the ‘LN’ column), the predictions in $\mathcal{A}_{\text{OOD-out}}(s)$ become noticeably lower. This effect becomes more apparent when the $x_{1}$ and $x_{2}$ axes are extended to the range $[-5,5]$ (see the ‘LN (Wider Range)’ column). With LN and a higher reward scale, the Q-values in $\mathcal{A}_{\text{OOD-out}}(s)$ remain closer to zero overall, making them appear relatively lower than those in $\mathcal{A}_{\mathcal{D}}$. Note that reward scaling is effective when used together with LN, as expected.
Figure 6: The normalized NTK map between $p=(0.32,-0.38)$ and $\mathcal{A}_{F}$ for $c_{\text{reward}}=1,100$.
To investigate this further, we plot the normalized NTK map to analyze the gradient similarity between the data boundary point $p=(0.32,-0.38)$ and the entire feasible action region $\mathcal{A}_{F}$. As shown in Figure 6, when $c_{\text{reward}}=1$, high gradient similarity is observed between $p$ and $\mathcal{A}_{\text{OOD-out}}(s)$. In contrast, as $c_{\text{reward}}$ increases to 100, the similarity decreases, indicating that the network perceives the OOD data as less similar to the training data. Consequently, the influence of positive gradient updates from the ID region ($x_{1}^{2}+x_{2}^{2}\leq 0.5^{2}$) on the OOD region is reduced during training, naturally leading to downward extrapolation.
@@ 4.3 PARS Algorithm
antmaze-u 78.6 87.5 74.0 0.0 0.0 27.5 97.9 93.5 97.0 97.8 93.8±$\pm 2.1$
antmaze-u-d 71.4 62.2 84.0 0.0 0.0 0.0 79.3 40.7 66.0 88.3 89.9$\pm$7.5
antmaze-m-p 10.6 71.2 61.2 0.0 0.0 0.0 85.9 74.7 38.5 84.0 91.2$\pm$3.9
antmaze-m-d 3.0 70.0 53.7 0.0 0.0 0.0 84.6 79.1 74.7 76.3 92.0$\pm$2.2
antmaze-l-p 0.2 39.6 15.8 0.0 0.0 0.0 64.3 35.3 43.9 60.4 84.8$\pm$5.9
antmaze-l-d 0.0 47.5 14.9 0.0 0.0 0.0 71.2 36.3 45.7 54.4 83.2$\pm$5.6
antmaze-ultra-p 0.0 13.3 16.1 0.0 0.0 0.0 0.6 4.4 20.6 22.4 66.4$\pm$4.4
antmaze-ultra-d 0.0 14.2 6.5 0.0 0.0 0.0 1.0 12.0 10.5 0.8 51.4$\pm$11.6
@@
pen-cloned 61.5 77.2 39.2 64.1 68.2 35.3 65.6 15.2 2.5 91.8 107.5$\pm$15.8
pen-expert 146.0 133.6 107.0 87.1 -1.5 121.2 119.9 117.3 45.4 154.1 152.7$\pm$1.0
door-cloned 0.1 0.8 0.4 -0.3 9.6 0.2 1.1 0.0 0.2 1.1 4.3$\pm$6.1
door-expert 84.6 105.3 101.5 -0.3 106.3 73.0 83.3 0.2 73.6 104.6 106.0$\pm$0.2
hammer-cloned 0.8 1.1 2.1 0.2 0.3 5.2 0.5 2.5 0.1 6.7 23.3$\pm$20.8
hammer-expert 117.0 129.6 86.7 25.1 28.5 75.9 103.3 86.6 24.8 133.8 133.5$\pm$0.4
relocate-cloned -0.1 0.2 -0.1 0.0 0.0 -0.1 0.0 -0.1 0.0 0.9 1.2$\pm$0.7
relocate-expert 107.3 106.5 95.0 -0.3 71.9 82.5 59.3 0.0 3.4 106.6 110.5$\pm$1.5
@@
halfcheetah-r 11.0 13.1 17.5 28.0 28.4 28.5 27.2 23.8 29.0 29.5 30.4$\pm$0.7
hopper-r 8.5 7.9 7.9 31.3 25.3 31.8 31.0 31.2 31.3 8.1 25.4$\pm$11.5
walker2d-r 1.6 5.4 5.1 21.7 16.6 17.0 2.2 5.3 21.5 18.4 21.8$\pm$0.2
halfcheetah-m 48.3 47.4 44.0 67.5 65.9 64.3 60.5 58.4 66.6 65.6 64.2$\pm$1.2
hopper-m 59.3 66.3 58.5 100.3 101.6 78.4 103.5 86.0 97.8 102.0 104.1$\pm$0.4
walker2d-m 83.7 78.3 72.5 87.9 92.5 91.0 92.4 86.4 91.6 82.5 97.3$\pm$2.5
halfcheetah-m-r 44.6 44.2 45.5 63.9 61.3 56.8 52.5 52.2 54.9 51.0 57.0$\pm$0.6
hopper-m-r 60.9 94.7 95.0 101.8 101.0 101.6 103.7 100.2 100.5 98.1 103.1$\pm$0.6
walker2d-m-r 81.8 73.9 77.2 78.7 87.1 91.3 95.6 91.6 88.7 77.3 95.8$\pm$1.4
halfcheetah-m-e 90.7 86.7 91.6 107.1 106.3 87.5 94.2 86.9 107.6 101.1 103.0$\pm$2.4
hopper-m-e 98.0 91.5 105.4 110.1 110.7 111.2 111.2 99.3 109.8 107.0 113.1$\pm$0.3
walker2d-m-e 110.1 109.6 108.8 116.7 114.7 114.2 109.3 112.0 105.0 111.6 111.8$\pm$0.7
@@
antmaze-u 99.0$\pm$0.6 98.4$\pm$1.9 95.2$\pm$1.6 99.4$\pm$0.8 90.1$\pm$10.8 99.4$\pm$0.9 99.7$\pm$0.8
antmaze-u-d 76.9$\pm$39.7 55.2$\pm$32.8 34.8$\pm$30.1 99.2$\pm$1.0 75.2$\pm$35.0 97.4$\pm$2.1 97.8$\pm$2.1
antmaze-m-p 94.4$\pm$3.0 91.2$\pm$3.8 83.4$\pm$2.3 97.4$\pm$1.4 95.1$\pm$6.3 96.8$\pm$1.9 99.1$\pm$1.8
antmaze-m-d 98.8$\pm$2.5 91.6$\pm$3.5 86.6$\pm$5.0 98.6$\pm$1.4 96.3$\pm$4.8 95.8$\pm$3.6 99.4$\pm$1.1
antmaze-l-p 87.3$\pm$5.6 60.4$\pm$21.5 56.0$\pm$3.9 93.0$\pm$2.5 75.0$\pm$14.7 71.4$\pm$30.9 96.2$\pm$3.0
antmaze-l-d 65.3$\pm$28.3 69.4$\pm$23.7 60.4$\pm$6.8 90.4$\pm$3.9 74.4$\pm$11.8 89.0$\pm$3.4 96.8$\pm$2.7
antmaze-ultra-p 21.3$\pm$19.0 0.0$\pm$0.0 13.3$\pm$5.8 8.8$\pm$16.5 6.9$\pm$2.7 0.0$\pm$0.0 86.5$\pm$4.4
antmaze-ultra-d 6.3$\pm$6.6 5.8$\pm$11.5 26.7$\pm$11.3 40.0$\pm$37.0 5.7$\pm$11.2 1.0$\pm$1.7 86.4$\pm$6.5
@@
pen-cloned -2.6$\pm$0.1 102.5$\pm$166.0 117.1$\pm$13.4 154.8$\pm$11.8 -1.6$\pm$1.6 134.1$\pm$7.2 155.4$\pm$3.1
door-cloned -$0$8.0$\pm$0.2 0.05$\pm$0.06 110.8$\pm$6.1 -0.34$\pm$0.0 53.3$\pm$35.3 102.1$\pm$26.8
hammer-cloned 0.24$\pm$0.03 -7.4$\pm$0.4 90.2$\pm$23.2 139.7$\pm$5.6 0.21$\pm$0.08 114.4$\pm$10.3 141.5$\pm$1.9
relocate-cloned -0.33$\pm$0.01 -1.5$\pm$0.5 -0.29$\pm$0.04 4.8$\pm$7.1 -0.34$\pm$0.01 1.5$\pm$1.1 53.8$\pm$7.7
@@
halfcheetah-r 26.5$\pm$3.4 92.7$\pm$5.7 60.9$\pm$5.0 91.5$\pm$2.5 32.9$\pm$8.1 6.8$\pm$3.9 100.1$\pm$2.9
hopper-r 10.0$\pm$1.5 95.3$\pm$9.2 48.5$\pm$38.9 90.2$\pm$19.1 17.7$\pm$26.0 12.4$\pm$1.8 109.7$\pm$5.3
walker2d-r 12.4$\pm$7.9 27.9$\pm$2.2 9.8$\pm$1.6 87.7$\pm$14.1 9.4$\pm$5.6 5.7$\pm$0.8 113.9$\pm$13.9
halfcheetah-m 78.9$\pm$1.3 103.3$\pm$1.4 70.4$\pm$2.3 95.5$\pm$1.5 77.0$\pm$2.2 56.6$\pm$0.8 107.0$\pm$5.0
hopper-m 100.9$\pm$0.6 106.3$\pm$1.7 86.2$\pm$26.3 91.4$\pm$27.8 100.7$\pm$0.8 104.8$\pm$2.6 111.5$\pm$0.4
walker2d-m 88.7$\pm$0.4 109.8$\pm$29.6 91.4$\pm$14.3 121.6$\pm$2.3 97.0$\pm$8.2 106.5$\pm$3.4 126.4$\pm$2.1
halfcheetah-m-r 50.3$\pm$28.3 95.6$\pm$1.7 55.4$\pm$5.1 90.1$\pm$1.3 62.1$\pm$1.1 53.2$\pm$5.4 98.5$\pm$1.0
hopper-m-r 103.9$\pm$1.8 101.7$\pm$14.8 95.3$\pm$7.2 78.9$\pm$24.5 101.4$\pm$2.1 103.4$\pm$6.6 107.0$\pm$1.4
walker2d-m-r 105.4$\pm$1.8 120.3$\pm$9.4 87.2$\pm$13.6 119.0$\pm$2.1 98.4$\pm$3.3 115.5$\pm$2.9 130.1$\pm$4.4
@@ 6.1 Experiment Setup
Hyperparameters.    We uniformly set $L_{I,i}=L_{I}$ and $U_{I,i}=U_{I}$ across all action dimensions $i$. We use $\Delta_{L}=|L_{I}|$ and $\Delta_{U}=|U_{I}|$, where $L_{I}$ and $U_{I}$ are scaled by either 100 or 1000 times the boundary of $\mathcal{A}_{F}$. Given that in the tasks considered, $L_{I}<0$ and $U_{I}>0$, we sample infeasible actions from the intervals $[2L_{I},L_{I}]$ and $[U_{I},2U_{I}]$, respectively. Furthermore, we enable the tuning of around 10 sets of hyperparameters, including $\alpha$ and TD3+BC’s $\beta$, similar to prior works (Wu et al., 2022; Nikulin et al., 2023; Tarasov et al., 2024). For AntMaze, strong performance can be achieved even with a single hyperparameter set, as shown in Appendix G. For detailed hyperparameters and implementation, please refer to Appendix F.3.
@@ C.1 Theoretical Justification of PARS
where $\mathrm{kNN}(s,a;\mathrm{ID})$ denotes the set of the k-nearest neighbors within $\mathrm{ID}$. We show that $T^{\pi}_{\mathrm{pars}}$ is a contraction under $|\cdot|_{\infty}$ norm:
@@ F.2 Baselines
halfcheetah-r 23.8$\pm$0.5
hopper-r 31.2$\pm$0.4
walker2d-r 5.3$\pm$9.3
pen-cloned 15.2$\pm$18.7
pen-expert 117.3$\pm$14.9
door-cloned 0.0$\pm$0.0
door-expert 0.2$\pm$0.0
hammer-cloned 2.5$\pm$3.2
hammer-expert 86.6$\pm$46.3
relocate-cloned -0.1$\pm$0.0
relocate-expert 0.0$\pm$0.0
@@
pen-cloned 65.6$\pm$18.8
pen-expert 119.9$\pm$11.2
door-cloned 1.1$\pm$1.6
door-expert 83.3$\pm$14.9
hammer-cloned 0.5$\pm$0.4
hammer-expert 103.3$\pm$16.2
relocate-cloned 0.0$\pm$0.0
relocate-expert 59.3$\pm$10.2
@@
antmaze-u 27.5$\pm$20.6
antmaze-u-d 0.0$\pm$0.0
antmaze-m-p 0.0$\pm$0.0
antmaze-m-d 0.0$\pm$0.0
antmaze-l-p 0.0$\pm$0.0
antmaze-l-d 0.0$\pm$0.0
pen-cloned 35.3$\pm$28.1
pen-expert 121.2$\pm$15.9
door-cloned 0.2$\pm$0.5
door-expert 73.0$\pm$2.2
hammer-cloned 5.2$\pm$6.3
hammer-expert 75.9$\pm$30.2
relocate-cloned -0.1$\pm$0.0
relocate-expert 82.5$\pm$7.2
@@
Goal-conditioned offline RL.    We obtain the HIQL scores from Park et al., 2024b, and the GC-IQL, WGCSL, DWSL, and GCPC scores from Zeng et al., 2024. For the umaze and umaze-diverse scores for HIQL, we run the experiments using the implementations provided for each respective algorithm in the official HIQL repository (https://github.com/seohongpark/HIQL). The resulting scores are as follows: umaze: 79.2$\pm$4.2, umaze-diverse: 86.2$\pm$5.7.
@@ Appendix G More Ablation Study
antmaze-umaze-diverse 91$\pm$3.0 58$\pm$2.1
antmaze-ultra-play 93$\pm$2.1 57$\pm$1.5
antmaze-ultra-diverse 94$\pm$1.8 54$\pm$2.3
@@
umaze 97.9 (MSG) 91.6 (GC-IQL) 93.8 93.3$\pm$3.2
umaze-diverse 88.3 (ReBRAC) 88.8 (GC-IQL) 89.9 84.7$\pm$5.7
medium-play 85.9 (MSG) 84.1 (HIQL) 91.2 89.4$\pm$2.9
medium-diverse 84.6 (MSG) 86.8 (HIQL) 92.0 86.2$\pm$5.6
large-play 64.3 (MSG) 86.1 (HIQL) 84.8 77.5$\pm$1.8
large-diverse 71.2 (MSG) 88.2 (HIQL) 83.2 83.5$\pm$3.6
ultra-play 22.4 (ReBRAC) 56.6 (GCPC) 66.4 60.2$\pm$7.6
ultra-diverse 14.2 (IQL) 54.6 (GCPC) 51.4 50.8$\pm$11.1
@@
umaze-diverse 71.6$\pm$10.5 89.9$\pm$7.5 88.6$\pm$5.9 87.5$\pm$6.9 82.3$\pm$8.2
medium-diverse 72.5$\pm$14.4 92.0$\pm$ 2.2 91.7$\pm$3.5 91.2$\pm$2.8 40.5$\pm$25.4
large-diverse 79.5$\pm$8.2 83.0$\pm$4.9 83.2$\pm$ 5.6 82.4$\pm$2.8 76.1$\pm$10.3
ultra-diverse 47.5$\pm$15.2 49.5$\pm$13.7 51.4$\pm$ 11.6 50.9$\pm$9.5 45.4$\pm$13.1

### v2

@@ 3.1 Extrapolation Error for OOD Actions Outside the Convex Hull of In-distribution Samples
To investigate this, we analyzed how the max-Q action, which maximizes the learned Q-function for a given state, differs from the dataset actions across various D4RL (Fu et al., 2020) datasets, similar to the approach of Yue et al. (2024). We trained $Q_{\phi}$ using observed state-action pairs from the dataset (offline SARSA; Sutton, 2018; Kumar et al., 2022) and then derived the max-Q policy using the loss function $\mathcal{L}_{\text{max-}Q}(\phi)=\mathbb{E}_{s\sim\mathcal{D}}\left[\text{max}_{a}Q_{\phi}(s,a)\right]$ .
@@ 4.1.1 Didactic Example
Figure 5: Results of training on a toy dataset using ReLU MLPs with vanilla regression (the ‘No LN’ column) and with LN (the ‘LN’ column), varying $c_{\text{reward}}$. The ‘LN (Wider Range)’ column shows results under the same conditions as ‘LN’, but with the $x_{1}$ and $x_{2}$ axes extended to the range [5,5][-5,5]. For more examples, please refer to Appendix D.
@@
As illustrated in the ‘No LN’ column, the Q-values exhibit catastrophic overestimation in $\mathcal{A}_{\text{OOD-out}}(s)$. Integrating LN helps mitigate the overestimation caused by linear extrapolation; however, $\mathcal{A}_{\text{OOD-out}}(s)$ remains overestimated relative to the in-distribution region. As the reward scale increases with LN applied (see the ‘LN’ column), the predictions in $\mathcal{A}_{\text{OOD-out}}(s)$ become noticeably lower. This effect becomes more apparent when the $x_{1}$ and $x_{2}$ axes are extended to the range [5,5][-5,5] (see the ‘LN (Wider Range)’ column). With LN and a higher reward scale, the Q-values in $\mathcal{A}_{\text{OOD-out}}(s)$ remain closer to zero overall, making them appear relatively lower than those in $\mathcal{A}_{\mathcal{D}}$. Note that reward scaling is effective when used together with LN, as expected.
Figure 6: The normalized NTK map between p=(0.32,0.38)p=(0.32,-0.38) and $\mathcal{A}_{F}$ for $c_{\text{reward}}=1,100$.
To investigate this further, we plot the normalized NTK map to analyze the gradient similarity between the data boundary point p=(0.32,0.38)p=(0.32,-0.38) and the entire feasible action region $\mathcal{A}_{F}$. As shown in Figure 6, when $c_{\text{reward}}=1$, high gradient similarity is observed between $p$ and $\mathcal{A}_{\text{OOD-out}}(s)$. In contrast, as $c_{\text{reward}}$ increases to 100, the similarity decreases, indicating that the network perceives the OOD data as less similar to the training data. Consequently, the influence of positive gradient updates from the ID region ($x_{1}^{2}+x_{2}^{2}\leq 0.5^{2}$) on the OOD region is reduced during training, naturally leading to downward extrapolation.
@@ 4.3 PARS Algorithm
[…]
@@
[…]
hammer-cloned 0.8 1.1 2.1 0.2 0.3 5.2 0.5 2.5 0.1 6.7 $3±\pm2$0.8
[…]
@@
[…]
hopper-m 59.3 66.3 58.5 100.3 101.6 78.4 103.5 86.0 97.8 102.0 $1±\pm0.4$
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 6.1 Experiment Setup
Hyperparameters.    We uniformly set $L_{I,i}=L_{I}$ and $U_{I,i}=U_{I}$ across all action dimensions $i$. We use $\Delta_{L}=|L_{I}|$ and $\Delta_{U}=|U_{I}|$, where $L_{I}$ and $U_{I}$ are scaled by either 100 or 1000 times the boundary of $\mathcal{A}_{F}$. Given that in the tasks considered, $L_{I}<0$ and $U_{I}>0$, we sample infeasible actions from the intervals $2L_{I},L_{I}]$ and $U_{I},2U_{I}]$, respectively. Furthermore, we enable the tuning of around 10 sets of hyperparameters, including $\alpha$ and TD3+BC’s $\beta$, similar to prior works (Wu et al., 2022; Nikulin et al., 2023; Tarasov et al., 2024). For AntMaze, strong performance can be achieved even with a single hyperparameter set, as shown in Appendix G. For detailed hyperparameters and implementation, please refer to Appendix F.3.
@@ C.1 Theoretical Justification of PARS
[…]
@@ F.2 Baselines
[…]
door-cloned $\pm0.0$
[…]
relocate-expert $\pm0.0$
@@
[…]
pen-expert $9±\pm11.$2
[…]
relocate-cloned $\pm0.0$
[…]
@@
[…]
antmaze-u-d $\pm0.0$
antmaze-m-p $\pm0.0$
antmaze-m-d $\pm0.0$
antmaze-l-p $\pm0.0$
antmaze-l-d $\pm0.0$
[…]
@@
[…]
@@ Appendix G More Ablation Study
[…]
@@
[…]
@@
[…]